A flat belt, 8 mm thick and 100 mm wide transmits power between two pulleys, running at 1600 m/min. The mass of the belt is 0.9 kg/m length. The angle of lap in the smaller pulley is 165° and the coefficient of friction between the belt and pulley is 0.3. If the maximum permissible stress in the belt is 2 MN/m2, find : 1. maximum power transmitted ; and 2. initial tension in the belt


1
Expert's answer
2021-06-04T04:54:02-0400

"\frac{T1}{
T2}= e^{μθ} = e^{0.3×2.88}= 2.372 … … . a"

"T_c = mv^2 = 0.9 \u00d7 (26.67)^2 = 640N"

"T = \u03c3 \u00d7 b \u00d7 t = T_c + T_1 \u2192 2 \u00d7 106 \u00d7 0.1 \u00d7 0.008 = 640 + T_1"

"T_1 = 1600 \u2212 640 \u2192 T_1 = 960N \u2026 \u2026 . b"

sub a in b ∶

"\frac{960}{
T2} = 2.372 → T_2 = 404.72N"

"P = (T_1 \u2212 T_2) \u00d7 v \u2192 P = (960 \u2212 404.72) \u00d7 26.67 = 14809.3W = 14.8kW"

"Tₒ =
\frac{T_1 + T_2 + 2T_c}{
2} =\frac{
960 + 404.72 + (2 × 640)}{
2 }= 1322.36N"

But when we use equation below , we get the same result which given for [Tₒ]

"Tₒ =
\frac{T_1 + T_2 + T_c}{
2} =
\frac{960 + 404.72 + 640}{
2} = 1002.36N"